• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 22:41:29 on October 30,1998

    Entry number 4215

    keyword=target NMR

    22:08 NMR Measurement

    UP: -14:18mV (fitted)
    DN: -14.24mV (fitted)
    ------->38.37% EPR
    ------->36.95% H20

    Note: Extract Pol. did not work on 1st attempt.
    After tweaking fit parameters was able to get convergence.
    Closer inspection showed that it was the resonance position
    that needed adjustment. It appears that the field is not quite
    right (off by ~1/2 G) after field rotation back to parrallel.
    -------------------------------------------------
    21:49 lasers off
    21:54/55 target motion (39246 ->20028)
    21:56 Rotation complete
    21:56 L1 on
    21:58 L2 on
    22:00 L3 on
    22:01 waiting for oven temp to rise (167->170)
    22:08 NMR/ RTD screenshot
    22:15 Target to beam position (039245)